y = (3x) / (x + 2).
Answer by math_tutor2020(3816)


It's not clear what your teacher wants here.
Please post the full instructions.

Since you posted this in the "inverses" section, I'll assume you want to find the inverse.
To do so, we will swap x and y, and then solve for y.
y = (3x)/(x+2)
x = (3y)/(y+2)
x(y+2) = 3y
xy+2x = 3y
2x = 3y-xy
2x = y(3-x)
(2x)/(3-x) = y
y = (2x)/(3-x)

Therefore, the inverse of y = (3x)/(x+2) is y = (2x)/(3-x)

y = (2x)/(3-x) is equivalent to y = (-2x)/(x-3) when we factor -1 out of the denominator.